OpenStudy (anonymous):

A 65.0 kg man sits in a 5.0 kg chair so that his weight is evenly distributed on the legs of the chair. Assume that each leg makes contact with the floor over a circular area with a radius of 1.15 cm. What is the pressure exerted on the floor by each leg?

OpenStudy (lastdaywork):

Can you calculate the total area of contact b/w the chair and the ground ??

OpenStudy (anonymous):

Must first calculate the total pressure on the legs 4 and the total pressure divided by 4. 65+5=70 70*9.8(gravity)=686(N) Area:3.14*(1.15*0.01)^2*4(4legs)=0.0016 686/0.0016=428750 428750/4=107187.5